OpenStudy (anonymous):

A 65kg person throw a 0.045kg snowball forward with a ground speed of 30m/s. A second person, with a mass of 60kg, catches the snowball. Both people on the skate. the first person is initially moving forward with speed of 2.5m/s, and the second person is initially at rest. what are the volocities of the two people after the snowball is exchanged? disregard friction between the skates and the ice

OpenStudy (anonymous):

u can help me ujjwal

OpenStudy (ujjwal):

yeah, i am trying to solve it for you.. This can solved by using 'conservation of momentum'.. i guess..

OpenStudy (anonymous):

yes, i confuse to solve the first person

OpenStudy (ujjwal):

for first person with snowball: initial momentum= (65+0.045)*2.5 final momentum= 65*x + 0.045*30 equating initial and final momentum you get, x=2.48 m/s So, final velocity of first person is 2.48 m/s can you solve for second person?

OpenStudy (anonymous):

why u use ini v=2.5, no v=2.5+30

OpenStudy (ujjwal):

because initially the snowball was in his hand and was moving at the speed at which he was moving..

OpenStudy (anonymous):

your answers is right, thanks u, i not sure

OpenStudy (ujjwal):

welcome buddy.. :)

OpenStudy (anonymous):

what is ground speed mean?

OpenStudy (ujjwal):

may be relative to ground